0:50 SX SOUL LUCI WFS powered up. Came up perfectly first try. Leaving warm until use in second half

4:20. LBTI will be handing over early so cooling OCAM

5:03 For real Barry sent the TTM Science field
  • AOref catalog mag 15.6 OCAM magnitude 17.3
  • Distance from teh moon 27deg
  • Optimize gain failed on tt the first time, safe skipping out on tt
  • Optimize gain fit to 0.05 for tt, but value probably shoudl have been close to 0.7-0.8 based on plots. The fit was poor. Manually bumped up
  • Optimize Gain plot for TTM target, poor fit:
    bin4_TT_Optimizegainplot.png
  • Gains used are :TT0.7, 0.19, 0.1
  • TN 20220608_051028, 051324, 054448
  • Seeing 0.64" on the luci2 wfs

6:00. Moving up to 60 degrees to reauthorize luci1 only for AO testing

Slewing to bin1 test target, , HD151670
  • Catalog magnitude R=9.0
  • Closing loop with NO GOPT to start the test
    • The pupils were very misaligned going from bin4 to bin1, center pupils was threatening to rip with high safeskips and driving the star off so interrupted command and manually centered
    • Camera lens is railed in X and pupil difference is bouncin between 0.09 and 0.35. We are proceeding with caution.
    • OCAM magnitude 10.0
    • tt 3.78 ho1 0.71 h02 0.60
    • Modal plot looks ok and does not suggest we are overdriving tt. Upper right and lower left pupils are darks and upper left and lower right are bright. Flux from double in 2 pupils and not the other two...
    • TN 062512 062543 062556
    • This is a double star, two stars of comparable magnitude with 0.29" angular separation.
6:27 Slewing to another bin1 BD+37 2879
  • Catalog magnitude R=8.3
  • Closing with NO GOPT
    • Camera lens well centered
    • OCAM magnitude 9.9
    • tt 3.48 ho1, 0.86 ho2 0.79, GOpt off, no NCPA. Bin1 900Hz, EM gain 533
    • Seeing 1.23"
    • TN 63701 , 63724
  • 6:50 reoptimizing, applying optical gain, NO NCPA
    • gains before Gopt 4.26 0.88 0.61
    • Optical gain applied 0.30, gains became 1.28 0.27 0.19
    • TN 06550, 070436
    • Optical gain is bouncing between 0.28 and 0.325, not too crazy. Really stabilised toward end of observation in particular.
  • 7:05 reoptimizing, gopt on, NCPA ON
    • Gains before optical gain applied 3/61, 0.91, 0.63
    • OPtical gain 0.289
    • NCPA on - visually verified NCPA trigger on in GOpt gui and NCPA coeficients populating, and can see it in the pupils
    • Seeing 0.95" DIMM
    • TN 070917, 071243
    • 7:14 starting to see lots of safe skips and degraded psf. TN 071436, 71735, 71749, 71804, 071917
    • up to 50% safeskips. Turning off NCPA but safeskips persist 72139, 72210, 72228
    • Optical gain is at 0.184 now.
    • Reoptimizing, but safeskips around the same level continue ->23%. Frame counter is incrementing. Pupils look ok
    • We opened loop before another tracking number taken this reset the high safeskip counter
  • 7:26 reoptimizing
    • Applying optical gain. 2.37 to start.
    • 7:32 Adsec arb offline 96% safeskips at the time -> focal station undefined, last command undefined when brought back online. Not associated with offset.
    • Returning_from_adsecArb_offline.png:
      Returning_from_adsecArb_offline.png
  • 7:37 Resending preset to BD+37
    • safeskipping during ramping of Ho2 during optimization
    • gains 5.04, 0.74, 0.48 before optical gain TN 074145
    • Apply Optical gain, optical gain value 0.317 to start TN 74256
    • NO NCPA
    • TN 074455, 074515, 074729
    • 7:47 Gopt seems to steadily be slowly dropping and is at 0.261 now.
    • 7:48 NCPA on Optical gain 0.277, 074942, 075002, 075015
    • Within2 minutes of starting NCPA Optical gain changed from 0.277 to 0.20
    • 7:51 loop opened
  • 7:52 Reclosing
    • gains before optical gain ->4.19, 0.18, 0.45 TN:075358
    • Optical gain -> 0.302
    • NCPA on. TN 075542, 75619, 75646
    • 7:58 Curled edge at 10:30 on OCAM. Looop opened. I wasn't fast enough for a screen shot or tracking number.
Tests have shown pretty conclusively that bin1
  • Gopt only was ok, but NCPA applied was not and was unstable
  • No NCPA still provides decent image quality, mostly trefoily...
  • Optimize gain for bin 1 - ranges may be too agressive. Safeskips of r ho2 although plot sdoes not look like it was obviously agressive
8:07 Sending TTM preset for DEC field. GS not found first try, so pointing check
  • Catalog Rmag 15.7
  • OCAM amgnitude 17.1 (17.2) (barry continues to est the faint end with an AO ref star on the other end of the patrol field from the science target...
  • bin4, freq 200, OCAM EM gain 600
  • Optimize gain fits are once again poor so just manually bumped up tt to something reasonable. Orginal fit had 0.05 for tt
    • Gains set to 0.7, 0.6, 0.1
  • 0.97 on the right WFS
  • Bumped down tt to 0.5...I think i might have gone a touch high. TN 083953
8:52 OCAM not live during closed loop - rebin recovered. TN085209

8:57 Resending preset
  • tt 0.44 0.45 0.1
  • OCAM mag mag still 17.2
  • TN 093445
  • some clouse, 28 photons per subapp... IQ on LUCI1 and LUCI2 comparable even with ESM
9:39. Clouds are becoming problematic so switching to MODS. Set safe state for SX WFS

-- JenniferPower - 08 Jun 2022

*
Topic revision: r3 - 08 Jun 2022, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback